OpenStudy (anonymous):

In a study of 100 common stocks, it was found that the annual dividend rates were normally distributed, with a mean of 4.0% and a standard deviation of 0.5% About how many of these stocks paid more than 5%?

OpenStudy (anonymous):

\[\frac{X-\mu}{\sigma}\] \[\frac{5-4}{.5}=2\] so you have a z score on the standard normal table of 2. You can look up the probability of being higher than that in a cart.

OpenStudy (anonymous):

chart*